# Large Language Models for Scientific Idea Generation: A Creativity-Centered Survey

### v1

@@ 3 Prompt-Driven Creativity: Steering LLMs Towards Novel Ideas
Prompt engineering is a powerful and efficient method to influence LLMs without costly retraining or fine-tuning. By designing precise instructions, role contexts, or structured templates within the input prompt, researchers can steer LLMs toward generating novel ideas and creative solutions (vatsal2024survey). As prompts function as external constraints and affordances that shape the model’s generative behavior, prompt driven methods align with the Press dimension of the 4P model trying to foster novelty while maintaining grounding in the model’s pretrained knowledge. Figure 1(b) schematically depicts these prompt-steering strategies and their placement within the idea-generation workflow. Building on this view, we discuss four main prompt-driven techniques for diversifying outputs: Persona & Role Priming, Constraint-Based & Adversarial Queries, Structured Creative Prompts, and Multilingual Prompting.
@@ 3.1 Persona & Role Priming
Assigning expert identities or roles to LLMs helps sharpen creative output. zhao2025assessingllms showed that role-based prompts—such as instructing the model to “act as a scientist”—can boost originality significantly on psychometric creativity tests. Supporting this, recent work in 2025 proposes frameworks that enhance LLM persona-awareness over multiple dialogue sessions, improving consistency and contextual relevance in outputs (liu-etal-2025-persona). Another study explores how persona conditioning influences model alignment and ethical decision-making, highlighting the importance of carefully designed persona prompts to guide LLM creativity while avoiding unintended biases (kim2025exploringpersonadependentllmalignment). Moreover, studies on human-AI co-creativity show that even ambiguous or counterintuitive personas can spur unexpected ideas, suggesting that nuanced role priming strategies can be highly effective in scientific ideation (wan2025usinggenerativeaipersonas).
@@ 3.2 Constraint-Based & Adversarial Queries
Prompting LLMs with constraints or adversarial instructions encourages them to explore less obvious, more innovative solutions. For example, asking models to challenge core scientific assumptions or forbidding common coding constructs can drive originality. Recent work introduces Denial Prompting, a method that incrementally imposes constraints on generated code to push models toward creative strategies (denial_prompting_2024). This approach is evaluated using a new creativity metric called NeoGauge, demonstrating that systematically denying routine solutions helps LLMs access novel areas of their generative space (denial_prompting_2024). These constraint-based methods effectively leverage the vast internal knowledge of LLMs while navigating the lower-probability, more inventive regions of output distributions.
@@ 3.3 Structured Creative Prompts
Structured prompts scaffold LLM reasoning, fostering more deliberate and diverse ideation. By guiding models through defined reasoning steps, they emulate human metacognitive strategies that enhance creativity. Chain-of-Thought (CoT) prompting, where models are encouraged to “think step by step,” improves reasoning accuracy and idea diversity (wei2022chainofthought). Building on this, recent advances like Consistency-Based Self-Adaptive Prompting (COSP) automate zero-shot prompt construction by leveraging the model’s own predictions and unlabeled data, enhancing performance on reasoning tasks without task-specific training (wan2023cosp).
In scientific domains, methods like Chain-of-Ideas (CoI), structures ideation by organizing literature into sequential knowledge chains that reflect the evolution of a research domain (li2024chain). Instead of exposing LLMs to unstructured data, CoI captures dependencies between past findings and emerging trends, enabling the model to identify gaps, trace idea development, and forecast future research directions. Likewise, Bit-Flip-Spark, structures idea generation into three stages: the Bit (baseline assumption), the Flip (challenging that assumption), and the Spark (novel insight that emerges) (o2025sparks). This scaffold systematically provokes cognitive conflict—an essential driver of creativity—by encouraging LLMs to question default premises and encourages the synthesis of ideas that transcend traditional disciplinary boundaries, a key step toward transformational creativity. Anticipated advances in LLMs, such as ChatGPT-5, are expected to amplify the benefits of structured prompting by enabling more detailed, collaborative reasoning and problem decomposition, further enriching creative outputs.
Scenario 2: When Prompting Alone Cannot Break Alignment Constraints Task: A researcher asks an LLM to “Think outside the box and propose a completely new programming paradigm beyond object-oriented or functional programming.” Aligned LLM response (current state). Despite being explicitly prompted to “be radically creative,” the aligned model produces only safe, incremental ideas, such as: A hybrid paradigm combining object-oriented and functional programming. Event-driven programming with AI-assisted code generation. Domain-specific languages for scientific computing. These outputs are plausible but unsurprising, reflecting the conservative bias from alignment training (e.g., RLHF). Unaligned / base model (hypothetical) response. A less-aligned model might generate more radical and unconventional concepts, for example: A self-modifying, biology-inspired code ecosystem where programs evolve like neural tissue. A programming language treating time as a first-class dimension These ideas may be impractical or incoherent, but they reflect true divergent exploration beyond the comfort zone of mainstream paradigms.
@@ 3.4 Multilingual Prompting
Recent studies reveal that multilingual prompting—translating prompts into multiple languages and aggregating the results—can significantly increase generative diversity and creativity. For instance, wang2025mlprompt found that cross-lingual prompting enables LLMs to draw on distinct linguistic and cultural priors, outperforming traditional diversity techniques such as temperature sampling. Similarly, a large-scale survey by vatsal2025survey across more than 250 languages highlights multilingual prompting as a broadly effective strategy for improving comprehension, reasoning, and creative output. The MLPrompt system further demonstrated how cross-lingual prompt reformulations can boost LLM comprehension and reasoning in complex tasks(wang2025mlprompt). Together, these findings suggest that linguistic diversity itself acts as a cognitive scaffold—encouraging LLMs to traverse alternative semantic frames and produce more original, cross-cultural ideas.
@@ 3.5 Takeaways and Limitations
Prompt-based techniques—such as persona prompting, constraint based queries, structured, and multilingual prompting—can effectively broaden LLM creativity by diversifying perspectives, reasoning paths, and linguistic representations. However, as shown in Scenario 3.3 even with strong creativity cues in the prompt—e.g., “Be radically original,” “Break all current conventions”— the aligned LLM remains trapped within a safe attractor basin created during RLHF and instruction tuning which prioritize safety and helpfulness over novelty. Mohammadi found that RLHF reduces output entropy and semantic variety, pushing models toward safer, less creative responses (mohammadi2024creativity). Similarly, explicit creativity cues can sometimes reduce fluency and flexibility (zhao2025assessingllms). This shows why simply prompting for creativity is insufficient: the model’s sampling distribution has been narrowed by alignment to avoid outputs that seem implausible or risky. Overcoming this requires inference-time scaling or multi-agent interaction, which actively expand the search space and allow for the emergence of more novel, cross-disciplinary ideas (wolf2023fundamental).
@@ 4 Search and sampling expansions: Inference-Time Exploration of Hypothesis Spaces
Inference-time scaling methods expand the capabilities of LLMs beyond single-pass decoding by enabling dynamic and iterative exploration without additional training. From a classical AI perspective (russell2016aima), these approaches can be framed as search problems: LLMs act as agents traversing vast combinatorial spaces of candidate hypotheses, guided by heuristics and evaluation functions.
By moving beyond random sampling, test-time scaling highlights the role of systematic search, where both the strategy of exploration and the abstraction level of representation shape the creative potential of generated ideas. The search strategy—ranging from focused local refinement to tree-structured branching—determines the breadth and depth of traversal, while the abstraction level—spanning from low-level outputs to higher-level conceptual hypotheses—defines the scope of novelty that can emerge. We expect that applying more systematic search at more abstract representational levels increases the likelihood of producing exploratory ideas, aligning this process with Boden’s notion of creativity as the structured traversal of a conceptual space (Boden2004).
@@ 4.1.1 Local Search: Single-Candidate Refinement
Sequential refinement methods iteratively improve a single candidate hypothesis, analogous to hill-climbing in classical AI (russell2016aima). This approach mirrors the scientific practice of proposing an initial idea and then refining it through feedback cycles.
For example, the Self-Refine method (liu2023selfrefine) uses internal feedback loops in which the LLM critiques its prior output and revises it without requiring additional training. PANEL (li2025PANEL) similarly decomposes reasoning into multiple steps, generating and critiquing each sequentially. ResearchAgent (baek2024researchagent) extends this paradigm by simulating peer review with multiple collaborating agents, though the refinement of each hypothesis remains sequential. CriticAL (li2024critical) integrates statistical hypothesis testing into this loop, enabling the detection of logical inconsistencies or unsupported claims.
A single-path search, while simple and low-cost, is fundamentally brittle. It may terminate with a suboptimal or uninteresting solution if the initial path leads to a local maximum, making it an incomplete method that risks missing a solution entirely. In contrast, beam search method designed to overcome this limitation by exploring multiple paths concurrently. Scenario 4.1.3 demonstrates how different search strategies address this trade-off between efficiency and exploration.
@@ 4.1.2 Population-Based Search: Beam-Style Refinement
Population-based methods maintain multiple candidate hypotheses simultaneously, propagating the most promising across iterations. This aligns with Local Beam Search, a natural extension of single-candidate local search (russell2016aima). Maintaining a population allows models to balance exploration and exploitation, reduce the risk of local optima, and improve diversity.
Modern LLM examples include MOOSE-Chem2 (yang2025moosechem), which run multiple instances or seeds in parallel, scoring candidates with internal or external evaluators. These approaches leverage ensemble diversity while maintaining refinement dynamics akin to local search.
Although local beam search reduces the brittleness of single-path refinement by maintaining a population of candidates, it remains constrained: once a branch is discarded, it cannot be recovered (Scenario 4.1.3). This means the method may still miss promising but initially under-performing directions. To address this limitation, researchers turn to tree-based approaches that allow a more systematic exploration of alternatives.
@@ 4.1.3 Tree Search: Branching Exploration
Branching exploration maps closely to tree-based informed search, where instead of tracking only a fixed-width beam, the search process can expand, backtrack, and revisit multiple paths, guided by heuristics that balance depth and breadth. This structured exploration mitigates premature pruning and supports the discovery of more diverse solutions. (russell2016aima).
@@
MC-NEST (liu2024mcnest) exemplifies a Monte Carlo Tree Search guided by Nash equilibrium principles, balancing exploration and exploitation through stochastic node selection strategies. Conceptually, this approach mirrors classical local search techniques such as stochastic hill climbing and simulated annealing (russell2016aima), in which the search tolerates occasional suboptimal moves to escape local minima. MAGIC (wang2025magic) applies multi-armed bandit strategies to allocate computational resources preferentially to promising branches. Monte Carlo Thought Search (kim2023montecarlothought) explores alternative synthesis pathways in catalysis design. Multi-agent systems, such as Virtual Scientists (johnson2024virsci) and the IRIS framework (feng-etal-2025-iris), implement branching exploration collaboratively, integrating both LLM and human feedback to prune and guide search.
Scenario 3: Search Mechanisms in Hallucination Mitigation Seed Idea: “Design a method to reduce llm factual hallucinations in medical questions.” 1. Local Search (Sequential Refinement): The model iteratively refines a single idea; Iteration 1: Add RAG with PubMed. Iteration 2: Integrate retrieval into every step of multi-hop reasoning. Iteration 3: Combine retrieval with a confidence estimator; rejects uncertain answers. Focuses on refining a single path, which may lock the model into one paradigm (e.g., RAG) and overlook other promising strategies like calibration or verification. 2. Local Beam Search (Parallel Refinement): Multiple candidate directions are explored simultaneously; RAG-based factual grounding -> iteration 1 -> iteration 2 -> … Auxiliary verifier to fact-checking -> iteration 1 -> iteration 2 -> … Uncertainty calibration methods -> iteration 1 -> iteration 2 -> … Training with hallucination-focused preference data -> iteration 1 -> … After each refinement round, weaker ideas (e.g., costly preference data) are dropped, while stronger ones continue. Balances multiple strategies before committing, but innovative yet noisy ideas may be pruned prematurely. 3. Tree Search (Branching Exploration): Structured hierarchical exploration of the idea space; Branch A: Retrieval-based approaches Hybrid BM25 + Dense retrievers Domain-specific knowledge bases Branch B: Verifier models External verifier LLM Ensemble fact-checking Branch C: Training-time alignment Hallucination-specific RLHF Adversarial training with hallucination prompts Branch D: Uncertainty-aware generation Selective abstention Calibrated probability outputs Heuristics like computational cost determine branch expansion, Preserve diverse strategies and enables backtracking, though strong heuristics are required to manage combinatorial growth.
@@
Across these mechanisms, we see a spectrum of trade-offs in how LLMs explore idea space during inference-time scaling. As shown in Scenario 4.1.3, single candidate refinement offers simplicity and efficiency but risks tunnel vision, converging too quickly on a narrow line of thought. Beam-Style refinement introduces diversity by allowing multiple candidates to coevolve, yet its pruning decisions are irreversible and may discard unconventional but valuable paths. Tree search provides the broadest and most systematic exploration, preserving multiple paradigms in parallel and enabling backtracking, though at significant computational cost.
@@ 4.2.1 Internal Self-Evaluation
A foundational feedback source is the model’s internal confidence, enabling self-critique and refinement without external input. Methods like Self-Refine (liu2023selfrefine) leverage the LLM’s ability to generate critical commentary on its own outputs and revise accordingly. PANEL (li2025PANEL) decomposes reasoning into discrete steps, each followed by model-generated critiques. CriticAL (li2024critical) further enhances this by applying statistical reasoning—such as hypothesis testing—to identify contradictions and weaknesses.
It is most effective when the model’s pretrained general linguistic knowledge suffices (e.g., logical consistency, surface fluency). However, numerous studies show that LLM self-confidence is often miscalibrated: models can be markedly overconfident in incorrect outputs or under-confident in correct ones, with accuracy–confidence alignment heavily influenced by prompt framing, distractors, and training objectives (geng2024surveyconfidence; leng2025taming; chhikara2025mindconfidencegapoverconfidence). As a result, reliance on internal heuristics risks bias reinforcement and even a form of reward hacking, where the model over-trusts its own scoring signals, strengthens spurious associations, or repeatedly exploits flaws in its confidence estimates without genuinely improving hypothesis quality.In Scenario 4.2.1, we compare the effects of internal confidence feedback versus peer-agent feedback on hypothesis evaluation.
Scenario 4: Internal Confidence vs Peer Agents in Hypothesis Evaluation Suppose an LLM proposes a novel compound, C22H28N6O4, as a potential inhibitor for a cancer-related enzyme. Internal Confidence Feedback. The model rates the compound as valid because its structure “looks” consistent with patterns learned during pretraining. It deems the structure plausible because it resembles valid chemical formulas and mimics motifs frequently seen in known inhibitors. Peer-Agent Feedback. A domain-specialized peer agent reviews the same compound and identifies structural instabilities or unrealistic bond angles that make it chemically unfeasible and flags it as unstable. Internal confidence is efficient and scalable, helping flag syntactically invalid outputs, but it may overestimate validity by focusing on surface-level coherence rather than underlying chemical feasibility. While peer-agent feedback enhances scientific rigor and domain awareness, though it requires additional computational and training resources to maintain reliable expert peers.

### v2

@@ 3 Prompt-Driven Creativity: Steering LLMs Towards Novel Ideas
Prompt engineering is a powerful and efficient method to influence LLMs without costly retraining or fine-tuning. By designing precise instructions, role contexts, or structured templates within the input prompt, researchers can steer LLMs toward generating novel ideas and creative solutions (Vatsal and Dubey, 2024). As prompts function as external constraints and affordances that shape the model’s generative behavior, prompt driven methods align with the Press dimension of the 4P model trying to foster novelty while maintaining grounding in the model’s pretrained knowledge. Figure 1(b) schematically depicts these prompt-steering strategies and their placement within the idea-generation workflow. Building on this view, we discuss four main prompt-driven techniques for diversifying outputs: Persona & Role Priming, Constraint-Based & Adversarial Queries, Structured Creative Prompts, and Multilingual Prompting.
@@ 3.1 Persona & Role Priming
Assigning expert identities or roles to LLMs helps sharpen creative output. Zhao et al. (2025) showed that role-based prompts—such as instructing the model to “act as a scientist”—can boost originality significantly on psychometric creativity tests. Supporting this, recent work in 2025 proposes frameworks that enhance LLM persona-awareness over multiple dialogue sessions, improving consistency and contextual relevance in outputs (Liu et al., 2025b). Another study explores how persona conditioning influences model alignment and ethical decision-making, highlighting the importance of carefully designed persona prompts to guide LLM creativity while avoiding unintended biases (Kim et al., 2025). Moreover, studies on human-AI co-creativity show that even ambiguous or counterintuitive personas can spur unexpected ideas, suggesting that nuanced role priming strategies can be highly effective in scientific ideation (Wan and Kalman, 2025).
@@ 3.2 Constraint-Based & Adversarial Queries
Prompting LLMs with constraints or adversarial instructions encourages them to explore less obvious, more innovative solutions. For example, asking models to challenge core scientific assumptions or forbidding common coding constructs can drive originality. Recent work introduces Denial Prompting, a method that incrementally imposes constraints on generated code to push models toward creative strategies (Lu et al., 2024b). This approach is evaluated using a new creativity metric called NeoGauge, demonstrating that systematically denying routine solutions helps LLMs access novel areas of their generative space (Lu et al., 2024b). These constraint-based methods effectively leverage the vast internal knowledge of LLMs while navigating the lower-probability, more inventive regions of output distributions.
@@ 3.3 Structured Creative Prompts
Structured prompts scaffold LLM reasoning, fostering more deliberate and diverse ideation. By guiding models through defined reasoning steps, they emulate human metacognitive strategies that enhance creativity. Chain-of-Thought (CoT) prompting, where models are encouraged to “think step by step,” improves reasoning accuracy and idea diversity (Wei et al., 2022). Building on this, recent advances like Consistency-Based Self-Adaptive Prompting (COSP) automate zero-shot prompt construction by leveraging the model’s own predictions and unlabeled data, enhancing performance on reasoning tasks without task-specific training (Wan et al., 2023).
In scientific domains, methods like Chain-of-Ideas (CoI), structures ideation by organizing literature into sequential knowledge chains that reflect the evolution of a research domain (Li et al., 2024a). Instead of exposing LLMs to unstructured data, CoI captures dependencies between past findings and emerging trends, enabling the model to identify gaps, trace idea development, and forecast future research directions. Likewise, Bit-Flip-Spark, structures idea generation into three stages: the Bit (baseline assumption), the Flip (challenging that assumption), and the Spark (novel insight that emerges) (O’Neill et al., 2025).
This scaffold is designed to induce cognitive conflict by prompting models to question default assumptions and adopt alternative premises, a mechanism commonly associated with creative ideation. Prior work shows that system-level instructions and structured reasoning scaffolds can reweight exploration and promote otherwise underrepresented trajectories within a fixed pretrained space (Reynolds and McDonell, 2021; Yao et al., 2023). In this more limited sense, structured prompting may be viewed as a procedural step toward transformational behavior, approximating aspects of conceptual space alteration through controlled search and assumption relaxation rather than true representational change.
Figure 4: Comparison between aligned and unaligned LLM responses, highlighting how alignment training suppresses divergent creative exploration.
@@ 3.4 Multilingual Prompting
Recent studies reveal that multilingual prompting—translating prompts into multiple languages and aggregating the results—can significantly increase generative diversity and creativity. For instance, Wang et al. (2025c) found that cross-lingual prompting enables LLMs to draw on distinct linguistic and cultural priors, outperforming traditional diversity techniques such as temperature sampling. Similarly, a large-scale survey by Vatsal et al. (2025) across more than 250 languages highlights multilingual prompting as a broadly effective strategy for improving comprehension, reasoning, and creative output. The MLPrompt system further demonstrated how cross-lingual prompt reformulations can boost LLM comprehension and reasoning in complex tasks(Wang et al., 2025c). Together, these findings suggest that linguistic diversity itself acts as a cognitive scaffold—encouraging LLMs to traverse alternative semantic frames and produce more original, cross-cultural ideas.
@@ 3.5 Takeaways and Limitations
Prompt-based techniques—such as persona prompting, constraint based queries, structured, and multilingual prompting—can effectively broaden LLM creativity by diversifying perspectives, reasoning paths, and linguistic representations. Empirical results from Zhao et al. (2025), using a TTCT-inspired framework with LLM-judge scoring aggregated over many samples, show that prompt-based techniques systematically shift creativity components. Interpreting originality as novelty and the mean of fluency, flexibility, and elaboration as valueness, instructional prompting increases novelty from 3.78 to 4.05 and valueness from 4.46 to 4.65. Chain-of-Thought prompting primarily boosts valueness (4.75 vs. 4.68) with negligible change in novelty (3.76 vs. 3.78). Persona-based prompting (e.g., natural scientist) achieve higher novelty (4.23) and valueness (4.62) than generic roles (4.1, 3.78). Overall, prompt-based methods enhance creativity by redistributing generation within existing knowledge, with limited impact on deeper exploratory novelty.
However, as shown in Scenario 4 even with strong creativity cues in the prompt—e.g., “Be radically original,” “Break all current conventions”— the aligned LLM remains trapped within a safe attractor basin created during RLHF and instruction tuning which prioritize safety and helpfulness over novelty. Some recent researches found that RLHF reduces output entropy and semantic variety, pushing models toward safer, less creative responses (Mohammadi, 2024; Kirk et al., 2023). Similarly, explicit creativity cues can sometimes reduce fluency and flexibility (Zhao et al., 2025). This shows why simply prompting for creativity is insufficient: the model’s sampling distribution has been narrowed by alignment to avoid outputs that seem implausible or risky (Wolf et al., 2023). To address these limitations, inference-time scaling and multi-agent interaction offer a promising direction by expanding the search space beyond one-pass prompting, potentially enabling more novel, cross-disciplinary ideas (Feng et al., 2025; Baek et al., 2024).
@@ 4 Search and sampling expansions: Inference-Time Exploration of Hypothesis Spaces
Inference-time scaling methods expand the capabilities of LLMs beyond single-pass decoding by enabling dynamic and iterative exploration without additional training. From a classical AI perspective (Russell and Norvig, 2020), these approaches can be framed as search problems: LLMs act as agents traversing vast combinatorial spaces of candidate hypotheses, guided by heuristics and evaluation functions.
By moving beyond random sampling, test-time scaling highlights the role of systematic search, where both the strategy of exploration and the abstraction level of representation shape the creative potential of generated ideas. The search strategy—ranging from focused local refinement to tree-structured branching—determines the breadth and depth of traversal, while the abstraction level—spanning from low-level outputs to higher-level conceptual hypotheses—defines the scope of novelty that can emerge. We expect that applying more systematic search at more abstract representational levels increases the likelihood of producing exploratory ideas, aligning this process with Boden’s notion of creativity as the structured traversal of a conceptual space (Boden, 2004).
@@ 4.1.1 Local Search: Single-Candidate Refinement
Sequential refinement methods iteratively improve a single candidate hypothesis, analogous to hill-climbing in classical AI (Russell and Norvig, 2020). This approach mirrors the scientific practice of proposing an initial idea and then refining it through feedback cycles.
For example, the Self-Refine method (Madaan et al., 2023) uses internal feedback loops in which the LLM critiques its prior output and revises it without requiring additional training. PANEL (Li et al., 2025d) similarly decomposes reasoning into multiple steps, generating and critiquing each sequentially. ResearchAgent (Baek et al., 2024) extends this paradigm by simulating peer review with multiple collaborating agents, though the refinement of each hypothesis remains sequential. CriticAL (Li et al., 2024b) integrates statistical hypothesis testing into this loop, enabling the detection of logical inconsistencies or unsupported claims.
A single-path search, while simple and low-cost, is fundamentally brittle. It may terminate with a suboptimal or uninteresting solution if the initial path leads to a local maximum, making it an incomplete method that risks missing a solution entirely. In contrast, beam search method designed to overcome this limitation by exploring multiple paths concurrently. Scenario 5 demonstrates how different search strategies address this trade-off between efficiency and exploration.
@@ 4.1.2 Population-Based Search: Beam-Style Refinement
Population-based methods maintain multiple candidate hypotheses simultaneously, propagating the most promising across iterations. This aligns with Local Beam Search, a natural extension of single-candidate local search (Russell and Norvig, 2020). Maintaining a population allows models to balance exploration and exploitation, reduce the risk of local optima, and improve diversity.
Modern LLM examples include MOOSE-Chem2 (Yang et al., 2025b), which run multiple instances or seeds in parallel, scoring candidates with internal or external evaluators. These approaches leverage ensemble diversity while maintaining refinement dynamics akin to local search.
Although local beam search reduces the brittleness of single-path refinement by maintaining a population of candidates, it remains constrained: once a branch is discarded, it cannot be recovered (Scenario 5). This means the method may still miss promising but initially under-performing directions. To address this limitation, researchers turn to tree-based approaches that allow a more systematic exploration of alternatives.
@@ 4.1.3 Tree Search: Branching Exploration
Branching exploration maps closely to tree-based informed search, where instead of tracking only a fixed-width beam, the search process can expand, backtrack, and revisit multiple paths, guided by heuristics that balance depth and breadth. This structured exploration mitigates premature pruning and supports the discovery of more diverse solutions. (Russell and Norvig, 2020).
@@
MC-NEST (Rabby et al., 2024) exemplifies a Monte Carlo Tree Search guided by Nash equilibrium principles, balancing exploration and exploitation through stochastic node selection strategies. Conceptually, this approach mirrors classical local search techniques such as stochastic hill climbing and simulated annealing (Russell and Norvig, 2020), in which the search tolerates occasional suboptimal moves to escape local minima. MAGIC (Xu et al., 2023a) applies multi-armed bandit strategies to allocate computational resources preferentially to promising branches. Monte Carlo Thought Search (Sprueill et al., 2023) explores alternative synthesis pathways in catalysis design. Multi-agent systems, such as Virtual Scientists (Johnson and Davis, 2024) and the IRIS framework (Feng et al., 2025), implement branching exploration collaboratively, integrating both LLM and human feedback to prune and guide search.
Figure 5: Comparison of sequential, parallel, and tree-based search mechanisms for the ideation task of reducing LLM hallucinations in medical question answering.
@@
Across these mechanisms, we see a spectrum of trade-offs in how LLMs explore idea space during inference-time scaling. As shown in Scenario 5, single candidate refinement offers simplicity and efficiency but risks tunnel vision, converging too quickly on a narrow line of thought. Beam-Style refinement introduces diversity by allowing multiple candidates to coevolve, yet its pruning decisions are irreversible and may discard unconventional but valuable paths (Zhou and Hansen, 2005; Meister et al., 2020; Cohen and Beck, 2019). Tree search provides the broadest and most systematic exploration, preserving multiple paradigms in parallel -though its ability to preserve unconventional hypotheses depends on the choice of heuristics- and enabling backtracking, a mechanism shown to be critical for recovering from premature commitments in recent LLM reasoning frameworks (Yao et al., 2023), albeit at significant computational cost.
@@ 4.2.1 Internal Self-Evaluation
A foundational feedback source is the model’s internal confidence, enabling self-critique and refinement without external input. Methods like Self-Refine (Madaan et al., 2023) leverage the LLM’s ability to generate critical commentary on its own outputs and revise accordingly. PANEL (Li et al., 2025d) decomposes reasoning into discrete steps, each followed by model-generated critiques. CriticAL (Li et al., 2024b) further enhances this by applying statistical reasoning—such as hypothesis testing—to identify contradictions and weaknesses.
It is most effective when the model’s pretrained general linguistic knowledge suffices (e.g., logical consistency, surface fluency). However, numerous studies show that LLM self-confidence is often miscalibrated: models can be markedly overconfident in incorrect outputs or under-confident in correct ones, with accuracy–confidence alignment heavily influenced by prompt framing, distractors, and training objectives (Geng et al., 2024; Leng et al., 2025; Chhikara, 2025). As a result, reliance on internal heuristics risks bias reinforcement and even a form of reward hacking, where the model over-trusts its own scoring signals, strengthens spurious associations, or repeatedly exploits flaws in its confidence estimates without genuinely improving hypothesis quality. n Scenario 6, we compare the effects of internal confidence feedback versus peer-agent feedback on hypothesis evaluation.
Figure 6: Contrasting internal confident and peer-based feedback in scientific hypothesis evaluation.